OpenStudy (anonymous):

Will give fan and medal! Simplify. square root of 9 over sixth root of 9 9 to the power of negative 2 over 3 9 to the power of negative 1 over 3 8 to the power of 3 over 10 9 to the power of 1 over 3

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@cool41 @rydertheepic @lovepurple

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Can you tell me how?

OpenStudy (anonymous):

You know what a square root is right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Well whats the square root of 9

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good now whats the square root of 9 over 6th root of 9

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

Well try to figure it out

OpenStudy (anonymous):

Look back in the book if you have one

OpenStudy (anonymous):

I dont have one its online

OpenStudy (anonymous):

k12?

OpenStudy (anonymous):

its gonna be 3/something

OpenStudy (anonymous):

campbell how did you do it you showed me yesterday

OpenStudy (anonymous):

Well look it up or Ill ask ab00t @abb0t GET IN HERE

OpenStudy (campbell_st):

well start by writing them using index notation \[\sqrt[n]{x} = x ^{\frac{1}{n}}\] so \[\sqrt{9} = 9^{\frac{1}{2}}\] and the other one is \[\sqrt[6]{9} = 9^{\frac{1}{6}}\] now the index law for division of the same base is subtract the powers \[9^{\frac{1}{2}} \div 9^{\frac{1}{6}} = 9^{\frac{1}{2} - \frac{1}{6}}\] and just finish it off for the answer

OpenStudy (anonymous):

9 1/4?

OpenStudy (anonymous):

well negative

OpenStudy (anonymous):

But thats not an answer choice.